what is the procedure to draw the graph of cothx , sechx.and cosechx ?


 

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vinod

Hyperbolic sine of x = sinh x = (ex - e-x)/2

Hyperbolic cosine of x = cosh x = (ex + e-x)/2

Hyperbolic tangent of x = tanh x = (ex - e-x)/(ex + e-x)

Hyperbolic cotangent of x = coth x = (ex + e-x)/(ex - e-x)

Hyperbolic secant of x = sech x = 2/(ex + e-x)

Hyperbolic cosecant of x = csch x = 2/(ex - e-x)

using above relation we can draw the graphs

 

GRAPHS OF HYPERBOLIC FUNCTIONS

y = sinh x
y = cosh x

 

y = tanh x
y = coth x

 

y = sech x
y = csch x

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..